OpenStudy (anonymous):

Which equation will have the same solution as 3 – 5(2x + 1) = 3(3 + x)? A. 11x = –13 B. 11x = 13 C. 13x = –11 D. 13x = 11

OpenStudy (nurali):

3 – 5(2x + 1) = 3(3 + x) 3-10x-5=9+3x -10x-2=9+3x 10x+3x=-2-9 13x=-11

OpenStudy (anonymous):

thanks

OpenStudy (nurali):

My Pleasure.